Al is ten years older than Jerry. At one point Al was three times Jerry's age. How old was Jerry when Al three times older than him?

Answer by ankor@dixie-net.com(22740)

Al is ten years older than Jerry.
a = j + 10
At one point Al was three times Jerry's age.
let y = "the point"
a - y = 3(j-y)
a = 3j - 3y + y
a = 3j - 2y
replace a with (j+10)
j + 10 = 3j - 2y
10 = 3j - j - 2y
10 = 2j - 2y
simplify, divide by 2
5 = j - y
We can write it:
y = j - 5
:
The problem with this problem is there are many solutions:
let a = 40 yrs old, then we know j = 30 yrs, find y
y = 30 - 5
y = 25 yrs ago Al(40) was 3 times Jerry (30)
check this
40 - 25 = 3(30-25)
15 = 3(5)
:
If a = 20, then j = 10, find y
y = 10 - 5
y = 5
20 - 5 = 3(10 - 5)
:
How about a = 27, then j = 17 and y = 12
27 - 12 = 3(17 - 12)